# The Exact Mixing Time for Trees of Fixed Order and Diameter

### v1

@@ 1 Introduction
For non-bipartite $G$, the distribution of $w_{t}$ converges to the stationary distribution $\pi$, where $\pi_{v}=\deg(v)/2|E|$. For bipartite $G$, we have the same convergence if we consider a lazy walk, in which we remain at the current state with probability $1/2$, at the cost of doubling the expected length of any walk. Herein, we will consider non-lazy walks on trees because they are simpler to analyze; our main result also holds for lazy walks, though we need to multiply our expressions by a factor of two. For an introduction to random walks on graphs, see Häggström [9] or Lovász [10].
@@
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+8d}{6(n-1)}$ if $\mbox{if }n-d\mbox{ is odd},$ (3)
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+11d-6}{6(n-1)}$ if nd is $\mbox{if $n-d$ is even}.$ (4)
@@ 2.1 The naive rule
We start by describing the naive rule from $v_{0}$ to $\pi$ on the path $P_{4}$ on vertices $(v_{0},v_{1},v_{2},v_{3})$. In the naive rule, we choose our target vertex up front (with the appropriate probability), and walk to that pre-chosen vertex. Our target probabilities are $(1/6,1/3,1/3,1/6)$ so the expected length of this naive stopping rule is
@@
This naive rule is optimal because $v_{3}$ is a halting state. Indeed, we get the same $19/6$ value from formula (6) via a different calculation.
Next, let’s consider a naive rule that is not optimal. This time, we follow the naive rule from leaf $w_{1}$ to $\pi$ on the star graph $S_{4}=K_{1,3}$ on vertices $\{w_{0},w_{1},w_{2},w_{3}\}$, where $w_{0}$ is the central vertex. Our target probabilities are $(1/2,1/6,1/6,1/6)$ and the expected length of the rule is
@@
This naive rule is not optimal because it does not have a halting state. For comparison, here is an optimal $(w_{1},\pi)$-stopping rule whose expected length is $3/2$: take one step, then take a second step with probability $1/2$.
@@ 3 The Mixing Time for a Double Broom
$\quad\mathsf{T}_{\mathsf{mix}}(G)$ $=\frac{2d^{3}+6d^{2}(\ell+r-2)+d(12\ell(r-2)-24r+37)+\ell(33-24r)+33r-42}{6(n-1)}.$
@@
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+8d}{6(n-1)}$ if nd is $\mbox{if $n-d$ is odd},$ (11)
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+11d-6}{6(n-1)}$ if nd is $\mbox{if $n-d$ is even}.$ (12)
@@
For a fixed order $n$ and diameters $13\leq d_{1}<d_{2}\leq n-$ 1, we have $\mathsf{T}_{\mathsf{mix}}(D_{n,d_{1}})<\mathsf{T}_{\mathsf{mix}}(D_{n,d_{2}})$.
@@
$\mathsf{T}_{\mathsf{mix}}(D_{n,d+1})-\mathsf{T}_{\mathsf{mix}}(D_{n,d})=\frac{-d^{2}+4d+n^{2}-2n-2}{2n-2}=\frac{n-1}{2}-\frac{d^{2}-2d+3}{2(n-1)}>0.$
@@ 4 Maximizing the Mixing Time
Starting from any tree $G\neq D_{n,d}$, we make incremental changes to the tree that increase the mixing time. This evolutionary process completes when we have created the balanced double broom $D_{n,d}$. Our evolution consists of three phases, as illustrated in Figure 2. In Phase One, we convert our tree into a caterpillar, and perhaps reduce the diameter to $s<d$. In Phase Two, we move leaves outwards in pairs, which results in either a double broom, or a double broom with one additional leaf located somewhere on the spine. In Phase Three, we move this additional leaf (if it exists), and balance the two sides of the double broom. Finally, if we reduced the diameter in phase one, we simply observe that $\mathsf{T}_{\mathsf{mix}}(D_{n,s})<\mathsf{T}_{\mathsf{mix}}(D_{n,d})$.
@@ 4.1 Phase One: Tree to Caterpillar
Let $G\in\mathcal{T}_{n,d}$ be a non-caterpillar with pessimal path $P=\{v_{0},v_{1},\ldots,v_{s}\}$, where $s\leq d$, and leaf $y\in V(G\backslash P)$ that is not adjacent to $P$. If $G^{*}=\sigma(G,P;y)$ then the following statements hold.
@@
    More generally, $H^{*}(v,v_{s})\leq H(v,v_{s})$ for any vertex $v\in V(G)$.
@@
    $H^{*}(\pi^{*},v_{s})<H(\pi,v_{s})$.
  4. (d)
    $\mathsf{T}_{\mathsf{mix}}(G)=H(v_{0},\pi)<H^{*}(v_{0},\pi^{*})\leq\mathsf{T}_{\mathsf{mix}}(G^{*}).$
@@
(a) By equation (9), it is clear that $H^{*}(v_{i},v_{s})=H(v_{i},v_{s})$ for all $0\leq i\leq s$, including for $v_{k}$.
(b) Moving $y$ closer to $v_{s}$ cannot increase hitting times to $v_{s}$, so $H^{*}(v,v_{s})\leq H(v,v_{s})$ for all $v\in V(G)$. In particular, we have $H^{*}(y,v_{s})=1+H^{*}(v_{k},v_{s})<1+H(x,v_{k})+H(v_{k},v_{s})=H(y,v_{s})$.
(c) Now let’s show that $H^{*}(\pi^{*},v_{s})<H(\pi,v_{s})$. When we change $G$ into $G^{*}$, we have $\deg^{*}(v_{k})=\deg(v_{k})+1$ and $\deg^{*}(x)=\deg(x)-1$, while all other degrees remain the same. As noted above, the hitting times to $v_{s}$ are non-increasing and the only $\pi_{i}$ that change are $\pi_{v_{k}}$ and $\pi_{x}$. Therefore,
$\quad H^{*}(\pi^{*},v_{s})-H(\pi,v_{s})$
$=\sum_{v}\pi^{*}_{v}H^{*}(v,v_{s})-\sum_{v}\pi_{v}H(v,v_{s})$
$\leq\pi^{*}_{v_{k}}H^{*}(v_{k},v_{s})+\pi^{*}_{x}H^{*}(x,v_{s})-\pi_{v_{k}}H(v_{k},v_{s})+\pi_{x}H(x,v_{s})$
$=\frac{1}{2(n-1)}\left(\deg^{*}(v_{k})H^{*}(v_{k},v_{s})+\deg^{*}(x)H^{*}(x,v_{s})-\deg(v_{k})H(v_{k},v_{s})-\deg(x)H(x,v_{s})\right)$
$<\frac{1}{2(n-1)}\left(H(v_{k},v_{s})-H(x,v_{s})\right)<0,$
where the first inequality follows from statement (b).
(d) Using statement (c), we have
@@
Let $P=\{v_{0},v_{1},\ldots,v_{s}\}$ be a pessimal path of $G$. By Lemma 4.5, we can transplant each leaf that is not adjacent to $P$, one at a time, so that each leaf is now adjacent to $P$. After each leaf transplant, $P$ is still the pessimal path due to (a) and (b). Meanwhile, the mixing time strictly increases by (d). We repeat this process until every leaf is adjacent to $P$. We now have a caterpillar $G^{*}\in\mathcal{T}_{n,s}$, where $s\leq d$, such that $\mathsf{T}_{\mathsf{mix}}(G)<\mathsf{T}_{\mathsf{mix}}(G^{*})$. ∎
@@ 4.2 Phase Two: Caterpillar to Double Broom or Near Double Broom
Consider a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$. For $0\leq k\leq d$, define the set $V_{k}$ to be the vertex $v_{k}$ and the leaves in $G\backslash P$ that are adjacent to it.
@@
Given a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$ that has leaves adjacent to $v_{i}$ and $v_{j}$ where $22\leq i\leq j\leq d-$ 2. Let $G^{*}=\tau(G,P;(i,i-1)\wedge(j,j+1))$, where $22\leq i\leq j\leq d-$ 2.
@@
Given a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$ and leaf $x$ adjacent to $v_{i}$ and leaf $y$ adjacent to $v_{j}$, where $22\leq i\leq j\leq d-$ 2. Let $G^{*}=\tau(G,P;(i,i-1)\wedge(j,j+1))$. Then $H^{*}(\pi^{*},v_{d})<H(\pi,v_{d})$.
@@
We show that
$02|E|(H^{*}(\pi^{*},v_{d})-H(\pi,v_{d}))=\sum_{v\in V}\deg^{*}(v)H^{*}(v,v_{d})-\sum_{v\in V}\deg(v)H(v,v_{d})<$ 0 .
@@
The overall change $H^{*}(\pi^{*},v_{d})-H(\pi,v_{d})$ is $1/2|E|$ times the sum of these eight terms. This sum is
@@
Given a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$ with leaf $x$ adjacent to $v_{i}$ and leaf $y$ adjacent to $v_{j}$, where $22\leq i\leq j\leq n-$ 2. Let $G^{*}=\tau(G,P;(i,i-1)\wedge(j,j+1))$. Then $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$.
@@
By Corollary 4.10, we can perform tree surgery $\tau(G,P;(i,i-1)\wedge(j,j+1))$ where $22\leq i\leq j\leq d-$ 2 to obtain $G^{*}$ with $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$. We can repeat this process as long as there are at least two leaves adjacent to $\{v_{2},v_{3},\ldots,v_{d-2}\}$, and the mixing time increases with each surgery. When we are done we either have a double broom, or a near double broom with one additional leaf adjacent to $\{v_{2},v_{3},\ldots,v_{d-2}\}$. ∎
@@ 4.3 Phase Three: Creating a Balanced Double Broom
Let $G\in\mathcal{T}_{n,d}$ be a near double broom with spine $v_{0},v_{1},\ldots,v_{d}$, where $\deg(v_{1})\geq\deg(v_{n-1})$, and with one more leaf $x$ incident with $v_{i}$ where $22\leq i\leq d-$ 2. Let $G^{*}$ be the double broom obtained via the leaf transplant $\tau(G,P;(i,d-1))$. Then $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$.
@@
The overall change $H^{*}(\pi^{*},v_{d})-H(\pi,v_{d})$ is $1/2|E|$ times the sum of these six terms. This sum is
$\quad-2(d-i-1)(H(v_{i},v_{i+1})-\deg(v_{i})-1)-H(v_{i},v_{d})-2(d-i-1)(\deg(v_{i})-1)$
$\qquad-H(v_{i},v_{d-1})-2(d-i-2)(d-i-1)+H(v_{d-1},v_{d})$
@@
So we can divide our desired inequality (16) by $2(d-i-1)$ to obtain the equivalent condition
@@
If $|L|\geq(n+1)/2$, then inequality (17) holds. Otherwise $|L|<(n+1)/2$, so we must have $i\leq d/2$ because the left end of the tree has $\ell+1>r$ leaves. So we can define
@@
corresponding to the middle and right end of the tree. We have $|M|=d-2i+1$ and $|R|<|L|$, and therefore
@@
Treat the double broom $G$ as having $\ell+1$ vertices adjacent to $v_{1}$ and $r$ vertices adjacent to $v_{d-1}$. Prior to equation (18), the argument is identical to the proof of Lemma 4.12, with $i=1$. Our desired condition (17) becomes $|E|\leq H(v_{1},v_{2})+d-4$. We know that $|R|\leq|L|-2$ and we have $|M|=d-2$. Therefore
@@
Assume the statement holds for trees (of all orders) with diameter at most $d-1$. Now consider a tree $G\in\mathcal{T}_{n,d}$ that is not the double broom $D_{n,d}$. If $G$ is not a caterpillar, then our first step is to replace $G$ with a caterpillar with a larger mixing time. Let $P=\{v_{0},v_{1},\ldots v_{s}\}$ be a pessimal path of $G$ where $s\leq d$. By Corollary 4.6, we can create a caterpillar $G^{*}\in\mathcal{T}_{n,s}$ with $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$.
@@
We are now in the endgame of our process. If our double broom is not balanced, we repeatedly apply Corollary 4.13 to create the balanced broom $B_{n,s}$, increasing the mixing time at each step. Finally, if $s<d$, then Lemma 3.5 shows that $\mathsf{T}_{\mathsf{mix}}(B_{n,s})<\mathsf{T}_{\mathsf{mix}}(B_{n,d})$.

### v2

# The Exact Mixing Time for Trees with Fixed Diameter

@@ 1 Introduction
For non-bipartite $G$, the distribution of $w_{t}$ converges to the stationary distribution $\pi$, where $\pi_{v}=\deg(v)/2|E|$. For bipartite $G$, we have the same convergence if we consider a lazy walk, in which we remain at the current state with probability 1/21/21 / 2, at the cost of doubling the expected length of any walk. Herein, we will consider non-lazy walks on trees because they are are simpler to analyze; our main result also holds for lazy walks, though we need to multiply our expressions by a factor of two. For an introduction to random walks on graphs, see Häggström [8] or Lovász [9].
@@
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+8d}{6(n-1)}$ $\mbox{if }n-d\mbox{ is odd},$ (3)
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+11d-6}{6(n-1)}$ $\mbox{if $n-d$ is even}.$ (4)
@@ 2.1 The naive rule
We start by describing the naive rule from $v_{0}$ to $\pi$ on the path $P_{4}$ on vertices $(v_{0},v_{1},v_{2},v_{3})$. In the naive rule, we choose our target vertex up front (with the appropriate probability), and walk to that pre-chosen vertex. Our target probabilities are (1/6,1/3,1/3,1/6)(1/6,1/3,1/3,1/6)( 1 / 6 , 1 / 3 , 1 / 3 , 1 / 6 ) so the expected length of this naive stopping rule is
@@
This naive rule is optimal because $v_{3}$ is a halting state. Indeed, we get the same 19/619/619 / 6 value from formula (6) via a different calculation.
Next, let’s consider a naive rule that is not optimal. This time, we follow the naive rule from leaf $w_{1}$ to $\pi$ on the star graph $S_{4}=K_{1,3}$ on vertices $\{w_{0},w_{1},w_{2},w_{3}\}$, where $w_{0}$ is the central vertex. Our target probabilities are (1/2,1/6,1/6,1/6)(1/2,1/6,1/6,1/6)( 1 / 2 , 1 / 6 , 1 / 6 , 1 / 6 ) and the expected length of the rule is
@@
This naive rule is not optimal because it does not have a halting state. For comparison, here is an optimal $(w_{1},\pi)$-stopping rule whose expected length is 3/23/23 / 2: take one step, then take a second step with probability 1/21/21 / 2.
@@ 3 The Mixing Time for a Double Broom
$\quad\mathsf{T}_{\mathsf{mix}}(G)$ $=\frac{1}{6(n-1)}\bigg{(}2d^{3}+6d^{2}(\ell+r-2)+d(12\ell(r-2)-24r+37)$
$\qquad+\ell(33-24r)+33r-42\bigg{)}.$
@@
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+8d}{6(n-1)}$ $\mbox{if $n-d$ is odd},$ (11)
$\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{(d-2)n-d+5}{2}-\frac{d^{3}-6d^{2}+11d-6}{6(n-1)}$ $\mbox{if $n-d$ is even}.$ (12)
@@
For a fixed order $n$ and diameters $3\leq d_{1}<d_{2}\leq n-1$, we have $\mathsf{T}_{\mathsf{mix}}(D_{n,d_{1}})<\mathsf{T}_{\mathsf{mix}}(D_{n,d_{2}})$.
@@
$\mathsf{T}_{\mathsf{mix}}(D_{n,d+1})-\mathsf{T}_{\mathsf{mix}}(D_{n,d})$ $=\frac{-d^{2}+4d+n^{2}-2n-2}{2n-2}$
$=\frac{n-1}{2}-\frac{d^{2}-2d+3}{2(n-1)}>0.$
@@ 4 Maximizing the Mixing Time
Starting from any tree $G\neq D_{n,d}$, we make incremental changes to the tree that increase the mixing time. This evolutionary process completes when we have created the balanced double broom $D_{n,d}$. Our evolution consists of three phases, as illustrated in Figure 2. In Phase One, we convert our tree into a caterpillar, and perhaps reduce the diameter to $s<d$. In Phase Two, we move leaves outwards in pairs, which results in either a double broom, or a double broom with one additional leaf located somewher on the spine. In Phase Three, we move this additional leaf (if it exists), and balance the two sides of the double broom. Finally, if we reduced the diameter in phase one, we simply observe that $\mathsf{T}_{\mathsf{mix}}(D_{n,s})<\mathsf{T}_{\mathsf{mix}}(D_{n,d})$.
@@ 4.1 Phase One: Tree to Caterpillar
Let $G\in\mathcal{T}_{n,d}$ be a non-caterpillar with leaf-to-leaf path $P=\{v_{0},v_{1},\ldots,v_{s}\}$, where $s\leq d$, and leaf $y\in V(G\backslash P)$ that is not adjacent to $P$. If $G^{*}=\sigma(G,P;y)$ then the following statements hold.
@@
    More generally, for $0\leq j\leq s$, we have $H^{*}(v,v_{j})\leq H(v,v_{j})$ for all $v\in V(G)$, with $H^{*}(y,v_{j})<H(y,v_{j})$, in particular.
@@
    $H^{*}(\pi^{*},v_{j})<H(\pi,v_{j})$ for $0\leq i\leq s$.
Proof.
(a) By equation (9), we have $H^{*}(v_{i},v_{j})=H(v_{i},v_{j})$ for all $0\leq i,j\leq s$.
(b) Moving $y$ closer to $v_{j}$ cannot increase hitting times to $v_{j}$, so $H^{*}(v,v_{j})\leq H(v,v_{j})$ for all $v\in V(G)$. In particular, we have
$H^{*}(y,v_{j})=1+H^{*}(v_{k},v_{j})<1+H(x,v_{k})+H(v_{k},v_{j})=H(y,v_{j}).$
(c) Now let’s show that $H^{*}(\pi^{*},v_{j})<H(\pi,v_{j})$. When we change $G$ into $G^{*}$, we have $\deg^{*}(v_{k})=\deg(v_{k})+1$ and $\deg^{*}(x)=\deg(x)-1$, while all other degrees remain the same. As noted above, the hitting times to $v_{j}$ are non-increasing and the only $\pi_{w}$ that change are $\pi_{v_{k}}$ and $\pi_{x}$. Therefore,
$\quad H^{*}(\pi^{*},v_{j})-H(\pi,v_{j})$
$=\sum_{v}\pi^{*}_{v}H^{*}(v,v_{j})-\sum_{v}\pi_{v}H(v,v_{j})$
$\leq\pi^{*}_{v_{k}}H^{*}(v_{k},v_{j})+\pi^{*}_{x}H^{*}(x,v_{j})-\pi_{v_{k}}H(v_{k},v_{j})+\pi_{x}H(x,v_{j})$
$=\frac{1}{2(n-1)}\bigg{(}\deg^{*}(v_{k})H^{*}(v_{k},v_{j})+\deg^{*}(x)H^{*}(x,v_{j})$
$\qquad\qquad-\deg(v_{k})H(v_{k},v_{j})-\deg(x)H(x,v_{j})\bigg{)}$
$\leq\frac{1}{2(n-1)}\left(H(v_{k},v_{j})-H(x,v_{j})\right)<0,$
where the first inequality follows from statement (b). ∎
Lemma 4.6.
Let $G\in\mathcal{T}_{n,d}$ be a non-caterpillar with pessimal path $P=\{v_{0},v_{1},\ldots,v_{s}\}$, where $s\leq d$, and leaf $y\in V(G\backslash P)$ that is not adjacent to $P$. If $G^{*}=\sigma(G,P;y)$ then $\mathsf{T}_{\mathsf{mix}}(G)=H(v_{0},\pi)<H^{*}(v_{0},\pi^{*})\leq\mathsf{T}_{\mathsf{mix}}(G^{*}).$
@@
Using Lemma 4.5 (c) for $j=s$, we have
@@
Let $P=\{v_{0},v_{1},\ldots,v_{s}\}$ be a pessimal path of $G$. We transplant each leaf that is not adjacent to $P$, one at a time, so that each leaf is now adjacent to $P$. After each leaf transplant, $P$ is still the pessimal path due to Lemma 4.5 (a) and (b). Meanwhile, the mixing time strictly increases by Lemma 4.6. We repeat this process until every leaf is adjacent to $P$. We now have a caterpillar $G^{*}\in\mathcal{T}_{n,s}$, where $s\leq d$, such that $\mathsf{T}_{\mathsf{mix}}(G)<\mathsf{T}_{\mathsf{mix}}(G^{*})$. ∎
@@ 4.2 Phase Two: Caterpillar to Double Broom or Near Double Broom
[…]
@@
Given a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$ that has leaves adjacent to $v_{i}$ and $v_{j}$ where $2\leq i\leq j\leq d-2$. Let $G^{*}=\tau(G,P;(i,i-1)\wedge(j,j+1))$, where $2\leq i\leq j\leq d-2$.
@@
Given a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$ and leaf $x$ adjacent to $v_{i}$ and leaf $y$ adjacent to $v_{j}$, where $2\leq i\leq j\leq d-2$. Let $G^{*}=\tau(G,P;(i,i-1)\wedge(j,j+1))$. Then $H^{*}(\pi^{*},v_{d})<H(\pi,v_{d})$ and $H^{*}(\pi^{*},v_{0})<H(\pi,v_{0})$.
@@
We prove the statement for $v_{d}$; the proof for $v_{0}$ follows by symmetry. We show that
$2|E|(H^{*}(\pi^{*},v_{d})-H(\pi,v_{d}))=\sum_{v\in V}\deg^{*}(v)H^{*}(v,v_{d})-\sum_{v\in V}\deg(v)H(v,v_{d})<0$
@@
[…]
@@
Given a caterpillar $G$ with spine $P=\{v_{0},v_{1},\ldots,v_{d}\}$ with leaf $x$ adjacent to $v_{i}$ and leaf $y$ adjacent to $v_{j}$, where $2\leq i\leq j\leq n-2$. Let $G^{*}=\tau(G,P;(i,i-1)\wedge(j,j+1))$. Then $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$.
@@
By Corollary 4.11, we can perform tree surgery $\tau(G,P;(i,i-1)\wedge(j,j+1))$ where $2\leq i\leq j\leq d-2$ to obtain $G^{*}$ with $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$. We can repeat this process as long as there are at least two leaves adjacent to $\{v_{2},v_{3},\ldots,v_{d-2}\}$, and the mixing time increases with each surgery. When we are done we either have a double broom, or a near double broom with one additional leaf adjacent to $\{v_{2},v_{3},\ldots,v_{d-2}\}$. ∎
@@ 4.3 Phase Three: Creating a Balanced Double Broom
Let $G\in\mathcal{T}_{n,d}$ be a near double broom with spine $v_{0},v_{1},\ldots,v_{d}$, where $\deg(v_{1})\geq\deg(v_{n-1})$, and with one more leaf $x$ incident with $v_{i}$ where $2\leq i\leq d-2$. Let $G^{*}$ be the double broom obtained via the leaf transplant $\tau(G,P;(i,d-1))$. Then $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$.
@@
[…]
$\quad-2(d-i-1)(H(v_{i},v_{i+1})-\deg(v_{i})-1)-H(v_{i},v_{d})$
$\qquad-2(d-i-1)(\deg(v_{i})-1)-H(v_{i},v_{d-1})$
$\qquad-2(d-i-2)(d-i-1)+H(v_{d-1},v_{d})$
@@
So we can divide our desired inquality (16) by $2(d-i-1)$ to obtain the equivalent condition
@@
[…]
@@
[…]
@@
[…]
@@
Assume the statement holds for trees (of all orders) with diameter at most $d-1$. Now consider a tree $G\in\mathcal{T}_{n,d}$ that is not the double broom $D_{n,d}$. If $G$ is not a caterpillar, then our first step is to replace $G$ with a caterpillar with a larger mixing time. Let $P=\{v_{0},v_{1},\ldots v_{s}\}$ be a pessimal path of $G$ where $s\leq d$. By Lemma 4.7, we can create a caterpillar $G^{*}\in\mathcal{T}_{n,s}$ with $\mathsf{T}_{\mathsf{mix}}(G^{*})>\mathsf{T}_{\mathsf{mix}}(G)$.
@@
We are now in the endgame of our process. If our double broom is not balanced, we repeatedly apply Corollary 4.14 to create the balanced broom $B_{n,r}$, increasing the mixing time at each step. Finally, if $s<d$, then Lemma 3.5 shows that $\mathsf{T}_{\mathsf{mix}}(B_{n,r})<\mathsf{T}_{\mathsf{mix}}(B_{n,d})$.